Dylan Jones kept careful records of the fuel efficiency of his new car. After the first eleven times he filled up the tank, he found the mean was 27.3 miles per gallon (mpg) with a sample standard deviation of 1.3 mpg.

  1. Compute the 95% confidence interval for his mpg.
Answer #1

Date Solution :- given that X = 27.3 mpg S = 1.3 mpg n = L degree of freedom = n-1= 11-1210 At 95% confidece level thet is RE